3gm of an oxide of metal is converted to chloride completly and it yeilded 5gm of chloride then eqivalent weight of metal is

Grade:12

1 Answers

Sunil Kumar FP
9 years ago
let the valency pf metal be x.
MOx(3gm)----MOy(5gm)
equibalent of mox=equivalent of moy
3/M/x+8 =5/M/x+35.5
calculating M/x=33.25

thanks and regards
sunil kr
askIItian faculty